### v5

@@ Abstract
We present SGS-SLAM, the first semantic visual SLAM system based on Gaussian Splatting. It incorporates appearance, geometry, and semantic features through multi-channel optimization, addressing the oversmoothing limitations of neural implicit SLAM systems in high-quality rendering, scene understanding, and object-level geometry. We introduce a unique semantic feature loss that effectively compensates for the shortcomings of traditional depth and color losses in object optimization. Through a semantic-guided keyframe selection strategy, we prevent erroneous reconstructions caused by cumulative errors. Extensive experiments demonstrate that SGS-SLAM delivers state-of-the-art performance in camera pose estimation, map reconstruction, precise semantic segmentation, and object-level geometric accuracy, while ensuring real-time rendering capabilities.
@@
SLAM 3D Reconstruction 3D Semantics 3D Segmentation
**footnotetext: These authors contributed equally to this workfootnotetext: Corresponding author  footnotetext: The demo video is available at https://youtu.be/y83yw1E-oUo.
@@ 1 Introduction
Dense Visual Simultaneous Localization and Mapping (SLAM) is a crucial problem in the field of computer vision. It aims to reconstruct a dense 3D map in an unseen environment while simultaneously tracking the camera poses in a real-time manner. Traditional visual SLAM systems [6, 29, 32, 27, 39, 9] stand out in sparse reconstruction using point clouds and voxels, but fall short in dense reconstruction. To extract dense geometric information for high-quality representation, learning-based SLAM methods [1, 36] have gained wild attention. They demonstrate proficiency in generating decent 3D global maps meanwhile exhibiting robustness on noises and outliers. In addition, drawing inspiration from the advancements in the neural radiance field (NeRF) [25], NeRF-based SLAM approaches [35, 47, 18, 44, 19, 37, 8] have made further progress. They excel in producing accurate and high-fidelity global reconstruction by capturing dense photometric information through differentiable rendering.
@@
Benefits from its rasterization of 3D primitives, Gaussian Splatting exhibits remarkably fast rendering speeds and allows direct gradient flow to each Gaussian’s parameters. This results in an almost linear projection between the dense photometric loss and parameters during optimization, unlike the hierarchical pixel sampling and indirect gradient flow through multiple non-linear layers seen in NeRF models. Moreover, the direct projection capability simplifies the addition of new parameters to the Gaussian field as separate channels, thereby enabling dynamic multi-channel rendering. Crucially, we integrate a semantic map into the 3D Gaussian field, essential for applications in robotics and mixed reality. This integration allows real-time rendering of appearance, depth, and semantic color.
@@
  • We introduce SGS-SLAM, the first semantic dense visual SLAM system grounded in 3D Gaussians. SGS-SLAM employs an explicit volumetric representation, enabling swift and real-time camera tracking and scene mapping. More importantly, it utilizes 2D semantic maps to learn 3D semantic representations expressed by Gaussians. Compared with previous NeRF-based methods which offer over-smooth object edges, SGS-SLAM provides high-fidelity reconstruction and optimal segmentation precision.
  • In SGS-SLAM, semantic maps provide additional supervision for optimizing parameters and selecting keyframes. We employ a multi-channel parameter optimization strategy where appearance, geometric, and semantic signals collectively contribute to camera tracking and scene reconstruction. Furthermore, SGS-SLAM utilizes these diverse channels for keyframe selection during the tracking phase, concentrating on actively recognizing objects seen earlier in the trajectory. This approach results in efficient and high-quality map reconstruction based on the chosen keyframes.
  • Utilizing semantic representation, SGS-SLAM provides a highly accurate disentangled object representation in 3D scenes, laying a solid foundation for downstream tasks such as scene editing and manipulation. SGS-SLAM facilitates the dynamic moving, rotating, or removal of objects in the map in real time. This is achieved by grouping Gaussians by specifying the semantic labels of the objects.
@@ 2.2 Neural Implicit SLAM
Methods based on NeRF [24], which handle complex topological structures and differentiable scene representation methods, have garnered significant attention, leading to the development of neural implicit SLAM methods [3, 21, 7, 20, 45]. iMAP [35] uses a single MLP for scene representation, which shows limitations in large-scale scenes. NICE-SLAM [47] uses pre-trained multiple MLPs for hierarchical scene representation. Co-SLAM [37] combines pixel set-based keyframe tracking with one-blob encoding. Go-SLAM [44] uses Droid-SLAM as a front-end tracking system and multi-resolution hash encoding [26] for mapping meanwhile integrated loop-closure detection and global optimization. However, these methods cannot utilize semantic information in maps. NIDS-SLAM [11] leverages the mature front-end tracking of ORB-SLAM3 [2] and Instant-NGP [26] for mapping but does not optimize joint semantic features for 3D reconstruction. DNS-SLAM [19] proposes a 2D semantic prior system that provides multi-view geometry constraints but does not optimize 3D reconstruction with semantic features. SNI-SLAM [46], a work parallel to ours, introduces semantic loss for geometric supervision but remains limited by the efficiency constraints of NeRF’s volume rendering.
@@
The outstanding performance and fast rasterization capabilities of 3D Gaussian Splatting [17] enable higher efficiency and accuracy on scene reconstruction. However, existing 3D-Gaussian-based SLAM systems [40, 16, 14] lack traditional effective tracking optimization, limiting tracking accuracy and the ability to recognize semantic information in scenes. We fuse semantic features into geometry and appearance and integrate 3D semantic features during keyframe selection in the tracking process. This allows us to obtain more effective and higher-resolution scene segmentation results while maintaining real-time performance.
@@ 3.1 Multi-Channel Gaussian Representation
where $K$ is the camera intrinsic matrix, $E_{t}$ is the extrinsic matrix capturing the rotation and translation of the camera at frame $t$, $l$ is the focal length. The influence of all Gaussians on this pixel can be combined by sorting the Gaussians in depth order and performing front-to-back volume rendering using the Max volume rendering formula [23]:
@@ 3.2.2 Keyframes Selection and Weighting
During the tracking phase of SLAM systems, keyframes are identified and stored simultaneously. These keyframes, providing different views of objects, are critical for mapping to refine 3D scene reconstruction. SGS-SLAM captures and stores keyframes at constant time intervals. Subsequently, keyframes associated with the current frame are chosen based on geometric and semantic constraints. Specifically, we randomly select pixels from the current frame and extract their corresponding Gaussians $G_{\rm sample}$ in the 3D scene. These Gaussians, $G_{\rm sample}$, are then projected onto the camera views of keyframes as $G_{\rm proj}$. The $G_{\rm proj}$ are evaluated based on the geometric overlap ratio:
@@ 3.2.3 Map Reconstruction
Compared to existing NeRF-based approaches [47, 15, 19, 46] that necessitate complex model architectures and feature fusion strategies, SGS-SLAM adopts explicit Gaussian representation for mapping. This approach allows for direct gradient flow to each parameter, resulting in high rendering speeds and optimal reconstruction quality. Compared to recent Gaussian-based methods [16, 40], SGS-SLAM incorporates geometric, appearance, and semantic features for multi-channel rendering. This enables the joint optimization of parameters across different channels, remarkably enhancing the efficiency and effectiveness of both mapping and segmentation processes.
@@ 4.1.2 Metrics
We use PSNR, Depth-L1 (on 2D depth maps), SSIM, and LPIPS to evaluate the reconstruction quality. For the evaluation of camera pose, we adopt the average absolute trajectory error (ATE RMSE). For semantic segmentation, we calculate mIoU score.
@@ 4.1.3 Baselines
We compare the tracking and mapping with state-of-the-art methods iMap [35], NICE-SLAM [47], Co-SLAM [37], ESLAM [15], and SplaTAM [16]. For semantic segmentation accuracy, we compare with NIDS-SLAM [11], DNS-SLAM [19], and SNI-SLAM [46].
@@ 4.2 Evaluation of Mapping and Localization
Figure 2: Qualitative comparison of our method and the baselines for reconstruction across three scenes from the Replica Dataset [34], with key details accentuated using colorful boxes. The results demonstrate that our method delivers more high-fidelity and robust reconstructions, particularly by capturing more detailed features of the objects in the scene.
@@
Table 2: Quantitative comparison in terms of Depth L1, ATE, memory usage, and FPS between our method and the baselines on the Replica dataset [34]. The values represent the average outcomes across eight scenes. The results of baselines are retrieved from [46]. Our method remarkably outperforms the baselines at Depth and ATE evaluations, and performs fairly on FPS metrics. Note that for parameter size, we only count the parameters of the model. SOTA performances are highlighted.
[…]
iMAP 4.645 3.118 4.153 9.92 2.23 1.82
@@ 4.4 Evaluation of Keyframe Optimization
Figure 3: The selected novel view evaluation of scene0000 from the ScanNet dataset [4]. The rendered views display the reconstructed objects such as bike, fridge, garbage bin, and guitar from novel views. Our method outperforms baselines by a large margin primarily due to the integration of keyframe optimization and semantic constraints. Note that the ground-truth for novel views is derived from the pre-established mesh provided by the ScanNet dataset.
@@ 4.6 Ablation Study
We perform the ablation of SGS-SLAM on the scene0000 of the ScanNet dataset [4] to evaluate the effectiveness of multi-channel feature supervision, and the keyframe optimization strategies.
Table 4: Ablation study of multi-channel optimization on the scene0000 of the ScanNet dataset [4]. The comparison involves settings where appearance, depth, and semantic supervision are removed.  means the metric is inapplicable.

### v6

@@ Abstract
We present SGS-SLAM, the first semantic visual SLAM system based on Gaussian Splatting. It incorporates appearance, geometry, and semantic features through multi-channel optimization, addressing the oversmoothing limitations of neural implicit SLAM systems in high-quality rendering, scene understanding, and object-level geometry. We introduce a unique semantic feature loss that effectively compensates for the shortcomings of traditional depth and color losses in object optimization. Through a semantic-guided keyframe selection strategy, we prevent erroneous reconstructions caused by cumulative errors. Extensive experiments demonstrate that SGS-SLAM delivers state-of-the-art performance in camera pose estimation, map reconstruction, precise semantic segmentation, and object-level geometric accuracy, while ensuring real-time rendering capabilities. The implementation code is available at https://github.com/ShuhongLL/SGS-SLAM.
@@
SLAM 3D Reconstruction 3D Segmentation
  footnotetext: The demo video is available at https://youtu.be/y83yw1E-oUo.
@@ 1 Introduction
Dense Visual Simultaneous Localization and Mapping (SLAM) is a crucial problem in the field of computer vision. It aims to reconstruct a dense 3D map in an unseen environment while simultaneously tracking the camera poses. Traditional visual SLAM systems [6, 32, 35, 30] stand out in sparse mapping using point clouds and voxels, but fall short in dense reconstruction. To extract dense geometric information for high-quality representation, learning-based SLAM methods [1, 39] have gained wild attention. They demonstrate proficiency in generating decent 3D global maps meanwhile exhibiting robustness on noises and outliers. Drawing inspiration from advancements in the neural radiance field (NeRF) [28], NeRF-based SLAM approaches [38, 51, 20, 48, 21, 41, 8] have made further progress. They excel in producing accurate and high-fidelity global reconstruction by capturing dense photometric information through differentiable rendering.
@@
Benefits from its rasterization of 3D primitives, Gaussian Splatting exhibits remarkably fast rendering speeds and allows direct gradient flow to each Gaussian’s parameters. This results in an almost linear projection between the dense photometric loss and parameters during optimization [18], unlike the hierarchical pixel sampling and indirect gradient flow through multiple non-linear layers seen in NeRF models. Moreover, the direct projection capability simplifies the addition of new channels to the Gaussian field, thereby enabling dynamic multi-channel feature rendering. Crucially, we integrate a semantic map into the 3D Gaussian field, essential for applications in robotics and mixed reality. This integration allows real-time rendering of appearance, depth, and semantic color.
@@
  • We introduce SGS-SLAM, the first semantic RGB-D SLAM system grounded in 3D Gaussians. SGS-SLAM employs an explicit volumetric representation, enabling swift and real-time camera tracking and scene mapping. More importantly, it utilizes 2D semantic maps to learn 3D semantic representations expressed by Gaussians. Compared with previous NeRF-based methods which offer over-smooth object edges, SGS-SLAM provides high-fidelity reconstruction and optimal segmentation precision.
  • In SGS-SLAM, semantic maps provide additional supervision for optimizing parameters and selecting keyframes. We employ a multi-channel parameter optimization strategy where appearance, geometric, and semantic signals collectively contribute to camera tracking and scene reconstruction. Furthermore, SGS-SLAM utilizes these diverse channels for keyframe selection during the tracking phase, concentrating on actively recognizing objects seen earlier in the trajectory.
  • Utilizing the semantic map, SGS-SLAM provides a highly accurate disentangled object representation in 3D scenes, laying a solid foundation for downstream tasks such as scene editing and manipulation. SGS-SLAM facilitates the dynamic moving, rotating, or removal of objects that is achieved by grouping Gaussians by specifying the semantic labels of the objects.
@@ 2.2 Neural Implicit SLAM
Methods based on NeRF [27], which handle complex topological structures and differentiable scene representation methods, have garnered significant attention, leading to the development of neural implicit SLAM methods [3, 11, 23, 22, 9, 49, 10]. iMAP [38] uses a single MLP for scene representation, which shows limitations in large-scale scenes. NICE-SLAM [51] uses pre-trained multiple MLPs for hierarchical scene representation. Co-SLAM [41] combines pixel-set-based keyframe tracking with one-blob encoding. Go-SLAM [48] uses Droid-SLAM [40] as the tracking system and multi-resolution hash encoding [29] for mapping. However, these methods cannot utilize semantic information in the map. NIDS-SLAM [13] leverages the tracking system of ORB-SLAM3 [2] and Instant-NGP [29] for mapping but does not optimize joint semantic features for 3D reconstruction. DNS-SLAM [21] proposes a 2D semantic prior system that provides multi-view geometry constraints but does not optimize 3D reconstruction with semantic features. DNS-SLAM [21] and SNI-SLAM [50] introduce semantic loss for geometric supervision but remain limited by the efficiency constraints of NeRF’s volume rendering.
@@
Utilizing the outstanding performance and fast rasterization capabilities of 3D Gaussian Splatting [19], Gaussian-based SLAM systems offer higher efficiency and accuracy on scene reconstruction [43, 18, 16, 26, 47, 7, 24]. However, existing Gaussian-based SLAM systems lack the ability to recognize semantic information in scenes. To bridge this gap, We utilize the semantic map during keyframe selection and integrate the semantic feature loss in the tracking and mapping process. This allows us to obtain more effective and higher-quality scene segmentation outcomes meanwhile preserving real-time processing performance.
@@ 3.1 Multi-Channel Gaussian Representation
where $K$ is the camera intrinsic matrix, $E_{t}$ is the extrinsic matrix capturing the rotation and translation of the camera at frame $t$, $l$ is the focal length. The influence of all Gaussians on this pixel can be combined by sorting the Gaussians in depth order and performing front-to-back volume rendering:
@@ 3.2.2 Keyframes Selection and Weighting
During the tracking phase of SLAM systems, keyframes are identified and stored simultaneously. These keyframes, providing different views of objects, are critical for mapping to refine 3D scene reconstruction. SGS-SLAM captures and stores keyframes at constant time intervals. Subsequently, keyframes associated with the current frame are chosen based on geometric and semantic constraints. Specifically, we randomly select pixels from the current frame and extract their corresponding Gaussians $G_{\rm sample}$ in the 3D scene. These sampled $G_{\rm sample}$ are then projected onto the camera views of keyframes as $G_{\rm proj}$, which are evaluated based on the geometric overlap ratio:
@@ 3.2.3 Map Reconstruction
Compared to existing NeRF-based approaches [51, 17, 21, 50] that necessitate complex model architectures and feature fusion strategies, SGS-SLAM adopts explicit Gaussian representation for mapping, resulting in high rendering speeds and optimal reconstruction quality. Compared to recent Gaussian-based methods [18, 43], SGS-SLAM incorporates geometric, appearance, and semantic features for multi-channel rendering. This enables the joint optimization of parameters across different channels, remarkably enhancing the efficiency and effectiveness of both mapping and segmentation processes.
@@ 4.1.2 Metrics
We use PSNR, Depth-L1, SSIM, and LPIPS to evaluate the reconstruction quality. To evaluate the camera pose, we use the average absolute trajectory error (ATE RMSE). For semantic segmentation, we calculate mIoU score.
@@ 4.1.3 Baselines
We compare the tracking and mapping with state-of-the-art methods NICE-SLAM [51], vMap [20], Co-SLAM [41], ESLAM [17], and SplaTAM [18]. For semantic segmentation accuracy, we compare with NIDS-SLAM [13], DNS-SLAM [21], and SNI-SLAM [50].
Figure 2: Qualitative comparison of our method and the baselines for reconstruction across three scenes from the Replica Dataset [37], with key details accentuated using colored boxes. The results demonstrate that our method delivers more high-fidelity and robust reconstructions.
@@ 4.2 Evaluation of Mapping and Localization
Table 2: Quantitative comparison in terms of Depth L1, ATE, and FPS between our method and the baselines on the Replica dataset [37]. The values represent the average outcomes across eight scenes. The FPS is evaluated by setting the same number of training iterations for all systems for fair comparison. The results of baselines are retrieved from [50]. Our method outperforms the baselines at Depth and ATE evaluations, and performs fairly on FPS metrics. SOTA performances are highlighted.
  Methods Depth L1 [cm]$\downarrow$ ATE Mean [cm]$\downarrow$ ATE RMSE [cm]$\downarrow$ Track. FPS [f/s]$\uparrow$ Map. FPS [f/s]$\uparrow$ SLAM FPS [f/s]$\uparrow$
@@ 4.4 Evaluation of Keyframe Optimization
Figure 3: The selected novel view synthesis of scene0000 from the ScanNet dataset [4]. The rendered views display the reconstructed objects such as bike, fridge, garbage bin, and guitar from novel views. Our method outperforms baselines by a large margin primarily due to the integration of keyframe optimization and semantic constraints. Note that the ground-truth for novel views is captured from the offline-reconstructed mesh provided by the ScanNet dataset.
@@ 4.6 Ablation Study
We perform the ablation of SGS-SLAM on the scene0000_00 of the ScanNet dataset [4] to evaluate the effectiveness of multi-channel feature supervision, and the keyframe optimization strategies.
Table 4: Ablation study of multi-channel optimization on the scene0000_00 of the ScanNet dataset [4]. The comparison involves settings where appearance, depth, and semantic supervision are removed.  means the metric is inapplicable.